# Unveiling Multiple Descents in Unsupervised Autoencoders

### v2

@@ 5 Double Descent in Nonlinear AEs
We demonstrate the double descent phenomenon using undercomplete FCN AEs (Figure LABEL:fig:mlp_ae_architecture), with additional results for CNNs in Appendix F.3, minimizing mean squared error (MSE) between inputs and outputs. Results are based on 5 to 15 random seeds, with implementation details in Appendix A. We trained models on contaminated datasets and tested on clean data, showing noise memorization (high test loss) versus signal learning (low test loss) and observed model, epoch, and sample-wise multiple descents. Train loss results are detailed in Appendix D, with additional findings on nonlinear synthetic datasets in Appendix F.4.
@@ 5.1 Model-Wise Double Descent
Figure LABEL:fig:non_linear_ae_heatmap illustrates double descent in nonlinear AEs, Concluding that nonlinear activations cause this phenomenon in unsupervised AEs. We define bottleneck-wise and hidden-wise double descent to distinguish between various model sizes and highlight the significance of our architectural choices. Undercomplete AEs ($m<D=50$) have a bottleneck set by the smallest layer, with the critical regime and second descent (which improves reconstruction) occurring when hidden layers exceed this size (above the white dashed line). Increasing the bottleneck shifts the critical regime. Hidden layers above the dashed line set the embedding layer as the bottleneck, while smaller hidden layers become the bottleneck (Figure LABEL:fig:mlp_ae_architecture). Overcomplete AEs ($m>D$) trivially learn the identity function and are excluded from this study (See results in Appendix F.2, Figure LABEL:fig:overcomplete_ae).
Sample and feature noise. Figure LABEL:fig:model_wise_dd_linear_subspace_test shows that increasing sample noise raises test loss, as higher noise levels dominate training and degrade performance. Higher noise levels also shape the double descent curve by shifting the critical regime to larger models, which are needed to memorize more noise. At low noise levels such as 0–20% sample noise, double descent is absent due to insufficient noise in the training data, making models in the critical regime learn the signal and preventing a test loss increase. This underscores noise as a key factor in the phenomenon. Figure LABEL:fig:model_wise_dd_single_cell_test demonstrates triple descent in single-cell RNA data, with similar test loss behavior across both interpolation phases. Appendix E provides additional evidence of double descent with feature noise, Figure LABEL:fig:model_wise_dd_feature_noise.
@@
SNR. Figure LABEL:fig:model_wise_dd_varying_snr_test (and LABEL:fig:model_wise_dd_varying_snr_feature_noise in the Appendix) show that SNR plays a key role in shaping test loss and revealing double descent. Negative SNR lets noise dominate the train set, causing memorization in the critical regime and exposing double descent. As SNR increases, noise influence diminishes, allowing better signal learning and lowering test loss. At SNR = 0 dB (Figure LABEL:fig:gaussian_model_wise_dd_varying_snr_test), double descent is absent because noise is not dominant enough to trigger memorization.
@@
Domain shift. Our experiments reveal non-monotonic behavior in test loss under domain shifts for the subspace data model (Figure 4). As the shift increases, test loss rises, but low shifts (0.1, 0.5) show no non-monotonicity due to strong source-target alignment, allowing interpolating models to perform well. Over-parameterized models further improve target reconstruction by reducing test loss. Subsection 6.1, Figure LABEL:fig:umap_single_cell_2, demonstrates double and triple descent in single-cell RNA data under real domain shift, reinforcing the phenomenon’s presence with real-world evidence.
@@
Anomalies. For the first time, we show that anomalies can cause double descent in test loss and non-monotonicity in anomaly detection (Cheng et al., 2021). Using data from 3.1, we analyze test loss and detection quality via ROC-AUC (Hanley & McNeil, 1982; Fawcett, 2006), which distinguishes clean (low reconstruction error) from anomalous samples (high error) (Malhotra et al., 2016; Borghesi et al., 2019; Lindenbaum et al., 2024). In the critical regime (Figure LABEL:fig:gaussian_sar_-15), anomaly memorization increases test loss for clean samples and reduces ROC-AUC. Larger models exhibit secondary descent, especially with low SAR and many anomalies, improving anomaly detection to match under-parameterized models while lowering test loss (Lerman & Maunu, 2018b, a; Han et al., 2022; Lindenbaum et al., 2024). Figure LABEL:fig:gaussian_sar_0 shows no double descent at high SAR, similar to previous results where contamination is not significant, yet critical regime models still perform the worst. Additional real-world insights are in Subsection 6.2.
@@
In conclusion, factors such as model nonlinearity, bottleneck size, and severe contaminations—like high noise levels, large domain shifts, many anomalies, and low SNR—affect double descent and increase test loss, occasionally shifting the critical regime.
@@
5.2 Epoch-Wise Double Descent
Inspired by the surprising results in (Nakkiran et al., 2021), which demonstrate epoch-wise double descent in a supervised setting, we study the existence of this phenomenon in our unsupervised framework. Figures LABEL:fig:epoch_wise_dd_sample_noise_varies_test and LABEL:fig:epoch_wise_dd_feature_noise_change (Appendix E) show how noisy samples (and features) affect the test loss respectively. Higher noise or lower SNR increases test loss, as shown in Figure LABEL:fig:epoch_wise_dd_snr_varies_test and Appendix E, Figure LABEL:fig:epoch_wise_dd_snr_varies_feature_noise. The phenomenon also occurs with domain shifts (Figure LABEL:fig:epoch_wise_domain_shift_test), where stronger shifts lead to higher test loss, shown in Figure LABEL:fig:epoch_wise_domain_shift_linear_subspace.
@@ 5.3 Sample-Wise Double Descent
We also analyze the effect of gradually increasing training samples while keeping model size fixed and observe non-monotonic trends in the test loss curve (Figures LABEL:fig:sample_noise_noise_per_sample_wise_double_descent, LABEL:fig:domain_shift_sample_wise_double_descent, and Appendix E Figure LABEL:fig:feature_noise_sample_wise_results). In some cases, this leads to double descent, as shown in Figure LABEL:fig:sample_noise_sample_wise_double_descent. Appendix D and Figure 15 present the training losses, along with additional results using more complex data models found in Appendices F.3 and F.4. These results confirm the effects of noise, SNR, and domain shift, which are consistent with the findings in Subsections 5.1 and 5.2.
@@ 6.1 Domain Adaptation
Domain shifts are common in machine learning, where differences between training and testing distributions can degrade performance on unseen data. Various domain adaptation methods (Peng et al., 2019; Chang et al., 2019; Zhou et al., 2022; Rozner et al., 2023; Yampolsky et al., 2023) aim to minimize this shift. In biology, this challenge, known as the ”batch effect,” arises when integrating datasets collected under different conditions (Tran et al., 2020).
This section explores the relationship between model size and its ability to address distribution shifts in real-world single-cell RNA data (Subsection 3.2). We demonstrate the benefits of over-parameterized models in unsupervised tasks under real domain shifts, revealing multiple descents. Related findings in supervised learning were shown by (Tripuraneni et al., 2021) and (Kausik et al., 2023) under different assumptions. Appendix F.1, Figure LABEL:fig:umap_source_target_single_cell visualizes source and target datasets using UMAP embeddings (McInnes et al., 2018). Figures LABEL:fig:domain_shift_test_losses, LABEL:fig:domain_shift_train_loss show test and train losses for models trained on source and tested on target datasets, where testing on the ’Wang’ dataset revealed triple descent.
We used KL-divergence (KLD) (Shlens, 2014a) to quantify the distribution shift between the source and target datasets, as shown in the test loss legend in Figure LABEL:fig:domain_shift_test_losses. Higher KLD values correspond to larger shifts and increased test loss, consistent with the simulated results in Figure 4.
@@ 6.2 Anomaly Detection
Unsupervised anomaly detection is vital in machine learning, with applications across various fields. AEs are widely used for this task (Chandola et al., 2009; Chen et al., 2018; Rozner et al., 2024; Lindenbaum et al., 2024). We trained AEs on both normal and anomaly data, detecting anomalies via reconstruction loss as detailed in Subsection 5.1.
@@ 7 Conclusions
Our study comprehensively studied double descent in unsupervised learning using AEs. Analytically and empirically, we showed that linear AEs lack double descent. In contrast, nonlinear AEs exhibit multiple descents and non-monotonic behaviors across model-wise, epoch-wise, and sample-wise levels for various data models and architectural designs. We analyzed the effects of sample and feature noise and emphasized the importance of bottleneck size in shaping double descent. Our findings also show that over-parameterized models enhance both reconstruction and downstream tasks like anomaly detection and domain adaptation, underscoring their real-world relevance. While the dataset sizes in our study may seem limited, conducting these experiments required significant computational resources, constraining our ability to run on massive datasets. Nevertheless, we argue that our evaluation remains extensive and conclusive. We hope this work inspires further research in other unsupervised learning frameworks and theoretical studies.
Impact Statement
This paper aims to contribute to the advancement of the field of Machine Learning. Our work focuses on refining the understanding of prior findings and does not have any potential societal implications, as it lacks direct applied outcomes.
Acknowledgments and Disclosure of Funding
The work of Tom Tirer is supported by the ISF grant No. 1940/23 and by MOST grant 0007091.
@@ Appendix A Implementation Details
This section provides complete implementation details for all experiments conducted in the paper. Illustrations of the subspace data model generation introduced in Subsection 3.1 for the scenarios of sample and feature noise, domain shift, and anomalies are displayed in Figures 9, 10, and 11 respectively.
Figure 9: Train samples generation for the scenarios of sample and feature noise with $p=0.5$. The first (leftmost) matrix depicts the latent vectors ${\boldsymbol{Z}}$. The second matrix illustrates the latent vectors being projected into a higher dimensional space, and the rightmost matrices contain clean (blue) and noisy (red) samples / features respectively. The test samples remain clean.
Figure 10: Train and test samples generation for the scenario of domain shift. The matrix on the left depicts the latent vectors ${\boldsymbol{Z}}$ and the two middle matrices represent the separation to source (${\boldsymbol{Z}}_{train}$) and target (${\boldsymbol{Z}}_{test}$). The two rightmost matrices illustrate the latent vectors of the train and test data being projected into a higher dimensional space with different matrices (${\boldsymbol{H}},{\boldsymbol{H}}^{{}^{\prime\prime}}$), resulting in a domain shift.
@@
Table 1: Parameters and hyper-parameters.
@@
For the subspace data model, nonlinear subspace data model, MNIST, and CIFAR-10 datasets, we generate 5000 samples for training and 10000 for testing across all scenarios (sample noise, feature noise, domain shift, and anomaly detection). Regarding the single-cell RNA data, we have focused on dataset number 4 from (Tran et al., 2020), which includes 5 distinct domains (biological batches) named ’Baron’, ’Mutaro,’ ’Segerstolpe,’ ’Wang,’ and ’Xin’, each representing 15 different cell types. Each cell (sample) in this dataset contains over 15000 genes (features). To facilitate the training of deep models while preserving the domain shift, we have retained the top 1000 prominent features. We utilize the ’Baron’ biological batch as our source data for the scenario of domain shift, comprising 5000 training samples, while the target batches are ’Mutaro’ (2122 samples), ’Segerstople’ (2127 samples), ’Wang’ (457 samples), and ’Xin’ (1492 samples). As for the sample and feature noise scenarios, we use the ’Baron’ domain for both sample and feature noise scenarios due to its largest sample size (8569). We allocate 5000 samples for training and introduce noise to specific samples and features, as described in subsection 3.1. The calculations of the SNR for both sample and feature noise cases are provided in Section B. The reserved 3569 samples are for testing. Please be aware that all the domains in this dataset are inherently noisy, reflecting their real-world nature. Therefore, even when no additional noise is applied ($p=0$), the data remains noisy. This may account for why the test loss does not decrease monotonically as the model size increases for cases with low noise levels, as shown in Appendix F.5, Figure LABEL:fig:sample_wise_final_ascent_cells.
@@
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}\|{\boldsymbol{y}}_{i}-\hat{{\boldsymbol{y}}}_{i}\|^{2}_{2}.$
Where $N$ is the number of training data samples, ${\boldsymbol{y}}_{i}$ is the true vector, and $\hat{{\boldsymbol{y}}}_{i}$ is the predicted vector. Due to contamination in the training dataset, the norm of train samples tends to be higher than that of the clean test samples. As the MSE loss is not scale-invariant, we opt to normalize both train and test losses only after the training process is complete, using $\frac{1}{N}\sum_{i=1}^{N}\|{\boldsymbol{y}}_{i}-\bar{{\boldsymbol{y}}}\|^{2}_{2}$, Where $\bar{{\boldsymbol{y}}}$ is the mean of $\{{\boldsymbol{y}}_{i}\}_{i=1}^{N}$. This strategy enables us to continue utilizing the MSE loss function while facilitating a fair and meaningful comparison between train and test losses.

### v3

@@ 6 Double Descent in Nonlinear AEs
We demonstrate the double descent phenomenon using FCN undercomplete AEs (Figure LABEL:fig:mlp_ae_architecture), minimizing mean squared error (MSE) between inputs and outputs. Results are based on 5 to 15 random seeds, with implementation details in Appendix A. We trained models on contaminated datasets, drawn from $\mathcal{P}_{{\bm{x}}}$ and tested on clean datasampled from $\mathcal{P}_{\tilde{{\bm{x}}}}$ to isolate the effect on the learned model, showing noise memorization (high test loss) versus signal learning (low test loss) and observed model, epoch, and sample-wise multiple descents. Unlike standard analyses of PCA (Josse & Husson, 2012; Minka, 2000), where both training and test data are typically noisy and the goal is to recover the underlying signal by attenuating noise-dominated components, our setup deliberately uses clean test data to isolate the effect of noise memorization in the learned model. Train loss results of the FCNs are detailed in Appendix D, with additional findings on nonlinear synthetic datasets in Appendix F.4. We also present double descent results for CNN AE models (Appendix A, Figure LABEL:fig:cnn_ae_architecture) in Subsection 6.1, Figures LABEL:fig:model_wise_sample_noise_mnist, LABEL:fig:model_wise_sample_noise_snr_mnist, LABEL:fig:model_wise_domain_shift_mnist_mnistm and in Appendix F.3.
@@ 6.1 Model-Wise Double Descent
Figure LABEL:fig:non_linear_ae_heatmap illustrates double descent in nonlinear AEs, concluding that nonlinear activations cause this phenomenon in unsupervised AEs. We define bottleneck-wise and hidden-wise double descent to distinguish between various model sizes and highlight the significance of our architectural choices. FCN undercomplete AEs ($m<D=50$) have a bottleneck set by the smallest layer, with the critical regime and second descent (improving reconstruction) occurring when hidden layers exceed this size (above the white dashed line). Increase in bottleneck shifts the critical regime. Hidden layers above the dashed line set the embedding layer as the bottleneck, while smaller hidden layers become the bottleneck (Figure LABEL:fig:mlp_ae_architecture). FCN overcomplete AEs ($m\geq D$) trivially learn the identity function and excluded from this study (see Appendix F.2, Figure LABEL:fig:overcomplete_ae).
Sample and feature noise. Figure LABEL:fig:model_wise_dd_linear_subspace_test (FCN) shows that increasing sample noise raises test loss, as higher noise levels dominate training and degrade performance. Higher noise levels also shape the double descent curve by shifting the critical regime to larger models, which are needed to memorize more noise. At low noise levels such as 0–20% sample noise, double descent is absent due to insufficient noise in the training data, making models in the critical regime learn the signal and preventing a test loss increase. This underscores noise as a key factor in the phenomenon. Figure LABEL:fig:model_wise_dd_single_cell_test (FCN) demonstrates triple descent in single-cell RNA data, with similar test loss behavior across both interpolation phases. Appendices E, F.3, F.4 provide additional evidence of double descent with feature noise. Figure LABEL:fig:model_wise_sample_noise_mnist presents double descent results for CNNs trained on the MNIST dataset. Additional results of CNNs trained on MNIST and CIFAR-10 datasets and further details regarding these experiments are presented in Appendix F.3.
@@
SNR. Figure LABEL:fig:model_wise_dd_varying_snr_test shows that SNR plays a key role in shaping test loss and revealing double descent. Negative SNR lets noise dominate the training set, causing memorization in the critical regime and exposing double descent. As SNR increases, noise influence diminishes, allowing better signal learning and lowering test loss. At SNR = 0 dB (Figures LABEL:fig:gaussian_model_wise_dd_varying_snr_test, LABEL:fig:model_wise_sample_noise_snr_mnist), double descent is absent or barely recognizable because noise is not dominant enough to trigger memorization.
@@
Domain shift. Our experiments reveal non-monotonic behavior in test loss under domain shifts for the subspace data model using FCNs (Figure LABEL:fig:model_wise_domain_shift_test). As the shift increases, test loss rises, but low shifts (0.1, 0.5) show no non-monotonicity due to strong source-target alignment, allowing interpolating models to perform well. Over-parameterized models further improve target reconstruction by reducing test loss. We also present double descent behavior when CNNs are trained on MNIST and tested on the MNIST-M dataset in Figure LABEL:fig:model_wise_domain_shift_mnist_mnistm. Subsection 7.1, Figure LABEL:fig:umap_single_cell_2, demonstrates double and triple descent in single-cell RNA data under real domain shift, reinforcing the phenomenon’s presence with real-world evidence.
@@
Anomalies. For the first time, we show that anomalies can cause double descent in test loss and non-monotonicity in anomaly detection (Cheng et al., 2021). Using data from 3.1, we analyze test loss and detection quality via receiver operating characteristic area under the curve (ROC-AUC) (Hanley & McNeil, 1982; Fawcett, 2006), which distinguishes clean (low reconstruction error) from anomalous samples (high error) (Malhotra et al., 2016; Borghesi et al., 2019; Lindenbaum et al., 2024). In the critical regime (Figure LABEL:fig:gaussian_sar_-15), anomaly memorization increases test loss for clean samples and reduces ROC-AUC. Larger models exhibit secondary descent, especially with low SAR and many anomalies, improving anomaly detection to match under-parameterized models while lowering test loss (Lerman & Maunu, 2018b; a; Han et al., 2022; Lindenbaum et al., 2024). Figure LABEL:fig:gaussian_sar_0 shows no double descent at high SAR, similar to previous results where contamination is not significant, yet critical regime models still perform the worst. Additional real-world insights are in Subsection 7.2.In this case we keep $\mathcal{P}_{{\bm{x}}}$ and $\mathcal{P}_{\tilde{{\bm{x}}}}$ the same to evaluate model performance on both clean and anomalous data to generate the ROC curve and compute the AUC, providing a meaningful metric for assessing anomaly detection and demonstrating the relevance of our findings to downstream tasks.
@@
In conclusion, factors such as model nonlinearity, bottleneck size, and severe contaminations, like high noise levels, large domain shifts, many anomalies, and low SNR, affect double descent and increase test loss, occasionally shifting the critical regime.
6.2 Epoch-Wise Double Descent
Inspired by the surprising results in (Nakkiran et al., 2021), which demonstrate epoch-wise double descent in a supervised setting, we study the existence of this phenomenon in our unsupervised framework. Figures LABEL:fig:epoch_wise_dd_sample_noise_varies_test and LABEL:fig:epoch_wise_dd_feature_noise_change (Appendix E) show how noisy samples (and features) affect the test loss, respectively. Higher noise or lower SNR increases test loss, as shown in Figure LABEL:fig:epoch_wise_dd_snr_varies_test and Appendix E, Figure LABEL:fig:epoch_wise_dd_snr_varies_feature_noise. The phenomenon also occurs with domain shifts (Figure LABEL:fig:epoch_wise_domain_shift_test), where stronger shifts lead to higher test loss, shown in Figure LABEL:fig:epoch_wise_domain_shift_linear_subspace.
@@ 6.3 Sample-Wise Double Descent
We analyze the effect of increasing training samples while keeping model size fixed and observe non-monotonic trends in the test loss curve (Figures LABEL:fig:sample_noise_noise_per_sample_wise_double_descent, LABEL:fig:domain_shift_sample_wise_double_descent, and Appendix E Figure LABEL:fig:feature_noise_sample_wise_results). In some cases, this leads to double descent, as shown in Figure LABEL:fig:sample_noise_sample_wise_double_descent. Appendix D and Figure 11 present the training losses, along with additional results using more complex data models found in Appendices F.3 and F.4. These results confirm the effects of noise, SNR, and domain shift, which are consistent with the findings in Subsections 6.1 and 6.2.
@@ 7.1 Domain Adaptation
Domain shifts are common in machine learning, where differences between training and testing distributions can degrade performance on unseen data. Various domain adaptation methods (Peng et al., 2019; Chang et al., 2019; Zhou et al., 2022; Rozner et al., 2023; Yampolsky et al., 2023) aim to minimize this shift. In biology, this challenge, known as the “batch effect”, arises when integrating datasets collected under different conditions (Tran et al., 2020).
This section explores the relationship between model size and its ability to address distribution shifts in real-world single-cell RNA data (Subsection 3.2). We demonstrate the benefits of over-parameterized models in unsupervised tasks under real domain shifts, revealing multiple descents. Related findings in supervised learning were shown by (Tripuraneni et al., 2021) and (Kausik et al., 2023) under different assumptions. Appendix F.1, Figure LABEL:fig:umap_source_target_single_cell visualizes source and target datasets using UMAP embeddings (McInnes et al., 2018). Figures LABEL:fig:domain_shift_test_losses, LABEL:fig:domain_shift_train_loss show test and train losses for models trained on source and tested on target datasets, where testing on the ‘Wang’ dataset revealed triple descent.
Since we are working with a real-world dataset, we do not have control over $\mathcal{P}_{{\bm{x}}}$ and $\mathcal{P}_{\tilde{{\bm{x}}}}$. In particular, we cannot control the distributional shift between the training and test sets. We used KL-divergence (KLD) (Shlens, 2014a) to quantify the distribution shift between the source and target datasets, as shown in the test loss legend in Figure LABEL:fig:domain_shift_test_losses. Higher KLD values correspond to larger shifts and increased test loss, consistent with the simulated results in Figure LABEL:fig:model_wise_domain_shift_test.
@@ 7.2 Anomaly Detection
Unsupervised anomaly detection is vital in machine learning, with applications across various fields. AEs are widely used for this task (Chandola et al., 2009; Chen et al., 2018; Rozner et al., 2024; Lindenbaum et al., 2024). We trained AEs on both normal and anomaly data, detecting anomalies via reconstruction loss as detailed in Subsection 6.1. As noted in Subsection 6.1 (the ‘Anomalies’ section), we align $\mathcal{P}_{{\bm{x}}}$ and $\mathcal{P}_{\tilde{{\bm{x}}}}$ to assess the model’s anomaly detection capabilities in the context of downstream tasks.
@@
8 Discussion
In this section, we identify and highlight several fundamental distinctions between double descent behavior in supervised learning and the unsupervised setting explored in this work.
Noise structure and data model. In supervised learning, the input $\bm{x}$} is typically clean, while noise is introduced in the labels $\bm{y}$}. In contrast, our AEs operate under an unsupervised reconstruction objective where the input serves as both source and target. Consequently, any corruption affects both sides of the learning signal (see last part of Section 2 and Section 3), inducing distinct generalization dynamics not observed in supervised settings.
Presence of double descent in linear models. Double descent has been extensively observed in supervised linear models such as linear regression (Belkin et al., 2019; Bartlett et al., 2020). However, our theoretical and empirical results demonstrate that linear unsupervised AEs do not exhibit double descent. This highlights a structural divergence between supervised and unsupervised learning, where nonlinearity becomes essential for the emergence of multiple descent phases.
Training objective vs. learning dynamics. While both supervised models and AEs often minimize the MSE, the training dynamics differ significantly. In supervised learning, the test distribution generally matches the training distribution. In contrast, our approach involves training on noisy inputs and evaluating on clean data, creating a distribution shift and affecting how generalization and overfitting emerge during training.
Interpretation of over-parameterization. In supervised tasks, interpolation is typically tied to the total number of parameters. Over-parameterization can be achieved by redistributing capacity across layers without necessarily increasing model size. However, in our setup, achieving interpolation requires joint scaling of the hidden and bottleneck layers (see Figure LABEL:fig:non_linear_ae_heatmap and Subsection 6.1).
@@ 9 Conclusions
Our study comprehensively studied double descent in unsupervised learning using AEs. Analytically and empirically, we showed that linear AEs lack double descent. In contrast, nonlinear AEs exhibit multiple descents and non-monotonic behaviors across model-wise, epoch-wise, and sample-wise levels for various data models and architectural designs. We analyzed the effects of sample and feature noise and emphasized the importance of bottleneck size in shaping double descent. Our findings also show that over-parameterized models enhance both reconstruction and downstream tasks like anomaly detection and domain adaptation, underscoring their real-world relevance. Although our experiments were conducted on datasets of moderate size due to computational constraints, we believe the scope and depth of the evaluation offer a strong and conclusive foundation. This work raises broader implications for the use of model capacity in unsupervised settings. It suggests that double descent is not unique to supervised learning and must be considered when designing AEs for real-world applications. Several open research questions emerge from our study, including the need for a deeper theoretical understanding of double descent in unsupervised models, how different levels of nonlinearity influence the phenomenon, and the role of architecture-specific factors in shaping generalization behavior. We hope these findings contribute to the growing understanding of learning dynamics in unsupervised models and inspire future research into both theoretical and practical aspects of model generalization in unsupervised learning frameworks.
Acknowledgment
TT was supported by ISF grant No. 1940/23 and MOST grant No. 0007091. OL was supported by MOST grant No. 0007341.
@@ Appendix A Implementation Details
This section provides complete implementation details for all experiments conducted in the paper. Illustrations of the subspace data model generation introduced in Subsection 3.1 for the scenarios of sample and feature noise, domain shift, and anomalies are displayed in Figures 7, 8, and 9 respectively.
Figure 7: Train samples generation for the scenarios of sample and feature noise with $p=0.5$. The first (leftmost) matrix depicts the latent vectors $\bm{Z}$}. The second matrix illustrates the latent vectors being projected into a higher dimensional space, and the rightmost matrices contain clean (blue) and noisy (red) samples / features respectively. The test samples remain clean.
Figure 8: Train and test samples generation for the scenario of domain shift. The matrix on the left depicts the latent vectors $\bm{Z}$} and the two middle matrices represent the separation to source (${\bm{Z}}_{train}$) and target (${\bm{Z}}_{test}$). The two rightmost matrices illustrate the latent vectors of the train and test data being projected into a higher dimensional space with different matrices (${\bm{H}},{\bm{H}}^{{}^{\prime\prime}}$), resulting in a domain shift.
@@
Table 1: Parameters and hyper-parameters
@@
For the subspace data model, nonlinear subspace data model, MNIST, and CIFAR-10 datasets, we generate 5000 samples for training and 10000 for testing across all scenarios (sample noise, feature noise, domain shift, and anomaly detection). Regarding the single-cell RNA data, we have focused on dataset number 4 from (Tran et al., 2020), which includes 5 distinct domains (biological batches) named ‘Baron’, ‘Mutaro,’ ‘Segerstolpe,’ ‘Wang,’ and ‘Xin’, each representing 15 different cell types. Each cell (sample) in this dataset contains over 15000 genes (features). To facilitate the training of deep models while preserving the domain shift, we have retained the top 1000 prominent features. We utilize the ‘Baron’ biological batch as our source data for the scenario of domain shift, comprising 5000 training samples, while the target batches are ‘Mutaro’ (2122 samples), ‘Segerstople’ (2127 samples), ‘Wang’ (457 samples), and ‘Xin’ (1492 samples). As for the sample and feature noise scenarios, we use the ‘Baron’ domain for both sample and feature noise scenarios due to its largest sample size (8569). We allocate 5000 samples for training and introduce noise to specific samples and features, as described in subsection 3.1. The calculations of the SNR for both sample and feature noise cases are provided in Section B. The reserved 3569 samples are for testing. Please be aware that all the domains in this dataset are inherently noisy, reflecting their real-world nature. Therefore, even when no additional noise is applied ($p=0$), the data remains noisy. This may account for why the test loss does not decrease monotonically as the model size increases for cases with low noise levels, as shown in Appendix F.5, Figure LABEL:fig:sample_wise_final_ascent_cells.
@@
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}\|{\bm{y}}_{i}-\hat{{\bm{y}}}_{i}\|^{2}_{2}.$
Where $N$ is the number of training data samples, ${\bm{y}}_{i}$ is the true vector, and $\hat{{\bm{y}}}_{i}$ is the predicted vector. Due to contamination in the training dataset, the norm of train samples tends to be higher than that of the clean test samples. As the MSE loss is not scale-invariant, we opt to normalize both train and test losses only after the training process is complete, using $\frac{1}{N}\sum_{i=1}^{N}\|{\bm{y}}_{i}-\bar{{\bm{y}}}\|^{2}_{2}$, Where $\bar{{\bm{y}}}$ is the mean of $\{{\bm{y}}_{i}\}_{i=1}^{N}$. This strategy enables us to continue utilizing the MSE loss function while facilitating a fair and meaningful comparison between train and test losses.